OpenStudy (anonymous):

The equation of a circle is (x + 12)2 + (y + 16)2 = (r1)2, and the circle passes through the origin. The equation of the circle then changes to (x – 30)2 + (y – 16)2 = (r2)2, and the circle still passes through the origin. What are the values of r1 and r2?

OpenStudy (anonymous):

A) r1 = 10 and r2 = 17 Eliminate B) r1 = 10 and r2 = 34 C) r1 = 20 and r2 = 17 D) r1 = 20 and r2 = 34

OpenStudy (anonymous):

sorry, the "eliminate" was not supposed to be there

OpenStudy (jdoe0001):

so... what's the center for the 1st equation of the circle?

OpenStudy (anonymous):

(-12, -16)?

OpenStudy (jdoe0001):

ok... so...what's the center for the 2nd equation of the circle?

OpenStudy (anonymous):

is it (30, 16)?

OpenStudy (anonymous):

what next?

OpenStudy (jdoe0001):

right they both pass through the origin, so we want to know their radius  so you have 2 points for each radius, the center of the circle, and the origin, to find the radius, or distance between then \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-12\quad ,&-16)\quad &(0\quad ,&0)\\ \quad \\ &(30\quad ,&16)\quad &(0\quad ,&0) \end{array}\qquad d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2} \)

OpenStudy (ranga):

An alternative solution: Both circles pass through the origin. Therefore, x = 0, y = 0 must satisfy the equations. Put (0,0) in the first equation and find r1 Put (0,0) in the second equation and find r2

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

Thank you both :) one more thing: how would you solve this? Find the equation of the parabola with focus (-4, 0) and directrix x = 4.

OpenStudy (anonymous):

sorry bout the "eliminate" thing again

OpenStudy (ranga):

One definition of a parabola is: The distance of any point on the parabola to the focus = distance of the point from the directrix: Let (x,y) represent a generic point on the parabola. The distance of the point (x, y) from (-4,0) = ? Distance of the point (x,y) from the line x = 4 is what? Equate them both to find the equation of the parabola.

OpenStudy (anonymous):

so would it be x^2=16y?

OpenStudy (ranga):

No. Try again.

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you :-)

OpenStudy (ranga):

you are welcome. :)